OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y

OpenStudy (anonymous):

(3, -1/2) (0, -2) (0, -11/2) (7, -5/2)

OpenStudy (anonymous):

use second equation.. x = 2 - 2y putting this in first, we obtain 4(2 - 2y) + 2y = 11 8 - 8y + 2y = 11 8 - 6y = 11 -6y = 3 y = -1/2 ...... (i) put this y in x = 2 - 2y u get, x = 2 - 2 ( -1/2 ) x = 2 + 2 ( 1/2 ) x = 2 + 1 x = 3. hence answer is ( 3, -1/2 )

OpenStudy (anonymous):

thanks!